OpenStudy (anonymous):

find the percent of change in altitdude of a weather ballon moves from 50ft to 95 ft describe the percent of change as an increase or decrease round to the nearest tenth

OpenStudy (amistre64):

find the difference in the heights then you can determine the percent of change

OpenStudy (amistre64):

50 * % = (95-50) 50 * n/100 = (95-50) solve for n

OpenStudy (anonymous):

thank you so much:)

OpenStudy (amistre64):

let me know how that works out :)

OpenStudy (anonymous):

i have one more question

OpenStudy (amistre64):

k

OpenStudy (anonymous):

let me post it